OpenStudy (anonymous):

Help?? The x-axis contains the base of an equilateral triangle RST. The origin is at S. Vertex T has coordinates (2h, 0) and the y-coordinate of R is g, with g > 0. Enter the coordinates for the midpoint of ST: (__,__) Enter the x- coordinate of R:__

OpenStudy (campbell_st):

the midpoint of ST is x = ( 0 + 2h)/2 and y = (0 +0)/2 midpoint is (h, 0) the midpoint will be the x coordinate of R R is the point (h, g)

OpenStudy (anonymous):

thank you so much;)